# AQM: A Refresh of the Abstract Qubit Model for Quantum Co-design

### v1

@@ IV AQM for Algorithm-Device codesign
States
Operation
Connectivity
Coherence
Readout
QA
$\checkmark$
Released
Critical to maintain high connectivity
$\checkmark$
Computational basis
QRW
Encoding methods can be different.
Released
Based on the problem, can be released.
$\checkmark$
Computational basis
MBQC
Physical qubits are not always alive.
Released (between physical qubits)
Resource state generation.
$\checkmark$
Arbitrary basis
AQS
$\checkmark$
Released
Based on the problem.
$\checkmark$
Computational basis (depending on problems).
QM
$\checkmark$
Released
Connectivity to computing component is necessary.
Long coherence time.
Released
IV.1 Applications without the complete AQM
@@
IV.1.1 Quantum Annealing
Analogous to the classical annealing method, quantum annealing is a QC paradigm to solve combinatorial optimization problems Das and Chakrabarti (2008); Hauke et al. (2020); Albash and Lidar (2018). Unlike the quantum circuit computing model, quantum annealing is an analog quantum computation framework. The optimization problem is encoded into the device Hamiltonian, whose ground state corresponds to the global optimum of the problem. Solving the problem corresponds to annealing the system states from an initial state to the ground state of the problem Hamiltonian.
@@ IV.1.1 Quantum Annealing
Therefore, quantum annealing does not require the support for Rule 2 of the AQM. As the computation is not driven by quantum gate operation, there is no need to support a full universal gate set. Instead, as long as the physical device can realize the Hamiltonian $H_{0}$ and $H_{1}$ that encodes the optimization problem, e.g., the nearest-neighbor spin-spin coupling and single-qubit Pauli-X rotations Johnson et al. (2011), and can slowly tune the system Hamiltonian and readout the final state of the system, the system can support the quantum annealing algorithm to solve the specific type of optimization problems.
For example, one type of physical system commonly considered in quantum annealing is the Ising spin glass, where $H_{0}$ contains single-site $\sigma^{(x)}$ terms, while $H_{1}$ contains two-site $Z$ couplings, i.e., $\sigma^{(z)}_{i}\sigma^{(z)}_{j}$. Although the coupling terms are restricted to two-body interactions, NP-hard problems can be cast into quadratic forms using ancilla Hauke et al. (2020); Lucas (2014). A survey of NP problems mapping to spin systems can be found in Ref. Lucas (2014). Note that realizing the Hamiltonian terms does not lead to universal control of individual spin in the system.
In order to keep the resource cost low when encoding the optimization problems into a quantum annealing setup, the physical systems need to have high connectivity, which is Rule 3 of the AQM. However, there are embedding techniques that allow mapping a quantum annealing problem onto physical systems with limited connectivity Djidjev et al. (2018); Zaribafiyan et al. (2017); Lechner et al. (2015).
In conclusion, the quantum annealing applications require Rule 1 and 4 of the complete AQM, while the Rule 2 and 5 can be released. In addition, the Rule 3 is also critical to keep the resource cost low.
IV.1.2 Quantum random walk.
Quantum random walk is a computation paradigm for QC, especially for solving graph problems. There are broad categories of quantum walk computations. In this section, we do not aim to give a thorough review of quantum random walk. For a review of quantum random walk for QC, we refer to Refs. Kadian et al. (2021); Venegas-Andraca (2012). Instead, we focus on the coined discrete-time quantum walk as an example.
@@
The quantum walk model can be used to implement universal QC Childs (2009); Singh et al. (2021). The quantum random walk model needs a different set of requirements compared to the AQM. For example, quantum information can be encoded not into the internal state of the qubit, but instead, into the spacial modes that can hold a physical particle Tang et al. (2018). Other than supporting a universal gate set on qubits, quantum walk computation only requires that the physical qubit can hop between different spatial modes that encode the quantum information to carry out the computation.
In summary, quantum walk applications require Rule 4 of the AQM. However, they have distinct requirements from the general AQM. Adaptable to problem demands, the quantum walk can operate on systems with more than two quantum states by encoding the coin system into a multi-level system based on walking choices, which differs from the Rule 1 of the AQM. Rule 2 of the AQM is released, as universal controllability and a universal gate set are not essential while controlled-walk operation suffices. Rule 3 can also be partially released, where the unique connectivity involves linking the systems encoding coin and space degrees of freedom (DOFs). Full connectivity is not required for unreachable space DOFs. For readout, the quantum walk model primarily emphasizes distribution in space DOFs Gong et al. (2021a); Esposito et al. (2022), making complete quantum information extraction (e.g., state tomography) unnecessary. Therefore, Rule 5 is also relaxed.
IV.1.3 Measurement-based quantum computation.
MBQC is a computation model different from the quantum circuit computation model Raussendorf and Briegel (2001); Raussendorf et al. (2003); Browne and Rudolph (2005); Briegel et al. (2009). In MBQC, the computation is driven by measurements of qubits in an entangled state of qubits. This entangled state is the resource of MBQC. A two-dimensional and higher-dimensional cluster state can be used as the resource state of MBQC Raussendorf et al. (2003); Van den Nest et al. (2006); Brown and Roberts (2020).
As demonstrated in Ref. Raussendorf et al. (2003), a set of universal gates can be equivalently performed in MBQC. Therefore, MBQC is equivalent to the quantum circuit computation model and it can support universal QC. However, compared to the requirements placed on the AQM, MBQC releases a few requirements. The physical qubits used in the MBQC are not accessible at all times, and they do not need to support a universal gate set for computation. Furthermore, the connectivity requirements are also different from the AQM, as it only requires that the physical qubits need to be able to prepare a specific resource state for MBQC. During the computation, connectivity between different physical qubits is not necessary.
@@
In summary, MBQC requires Rule 4 and 5 of the AOM. In contrast, MBQC relaxes Rule 1 of the AQM, where the physical qubits do not need to be alive throughout the computation. Rule 2 of the AQM is also released, as the physical qubits do not need to perform entangling gates. Rule 3 is supported in a different manner, which refers to the connectivity of the resource state.
IV.1.4 Analogue Quantum Simulation
@@
There are two main quantum simulation categories based on the type of quantum systems used. One involves digital quantum computers equipped with algorithms to compute physical properties, such as the VQE algorithm designed for determining molecular ground states and energies Cerezo et al. (2021). The other method focuses on employing quantum devices for analogue simulations Monroe et al. (2021); Lv et al. (2018); Ebadi et al. (2021); Daley et al. (2022), which is the emphasis of this subsection.
@@
Analogue quantum simulations can address low-energy steady states or dynamics of physical systems. Solving the ground state of a physical system can use adiabatic evolution, akin to the discussion of the quantum annealing method (see Sec. IV.1.1). Alternatively, when the problem directly maps to a quantum system, i.e., the quantum simulator can realize the target Hamiltonian, manipulating the simulator can simulate the dynamics of the target system. Therefore, the measurement of the quantum simulator can help to reveal the physical properties of the target problem. For example, the Ising model can naturally map to a neutral atom lattice Ebadi et al. (2021). The quantum many-body phase transition can be directly observed by manipulating the atom interactions Ebadi et al. (2021); Bloch et al. (2008).
@@
In summary, analogue quantum simulation still demands Rule 1 and 4 of the AQM. As the simulator does not need to support a universal gate set for computing, or achieving universal control on each qubit, as long as the target Hamiltonian can be realized, Rule 2 of the AQM is released. As the connectivity and the readout capability depend on the physical problems being simulated, Rule 3 and 5 of the AQM have the potential to be released.
IV.1.5 Quantum Memory
@@
Quantum memory is an application of quantum systems, which can preserve quantum information for an extended period of time. Unlike the other quantum computation paradigms and quantum simulation schemes, quantum memory does not aim to perform computation. Therefore, the quantum memory application has different demands on the AQM Liu et al. (2023b).
@@
To support quantum memory applications, Rule 1 and 4 of the AQM are required. The AQM Rule 2 can be released, as a quantum system does not need to support universal control of individual qubits and a universal gate set for quantum computation. Rule 5 can also be released as when the quantum memory acts as a computing module in the future QPU architecture design, its readout capability is not necessary. Rule 3 is differently demanded, as the connectivity of the qubits inside the quantum memory modules is not essential, while connectivity to computing qubits is necessary Liu et al. (2023b).
IV.2 Physical systems supporting the AQM
@@
In this section, we examine the physical system from the lower stacks of the AQM. We do not aim to provide a thorough review of useful physical systems for QC and QIP. Instead, we mainly focus on three types of physical systems: (1) physical systems with strong support of the AQM, e.g., superconducting circuits, trapped ions, and neutral atoms, (2) systems with unique strength and weakness in supporting the AQM, e.g., nitrogen-vacancy centers in diamond crystals, and photonic systems, (3) systems with incomplete support of the AQM, e.g, the quantum memory systems for photons.
IV.2.1 Physical systems for AQM support
@@
Superconducting circuits Devoret et al. (2004); Girvin et al. (2009); Kjaergaard et al. (2020); Kockum and Nori (2019); Blais et al. (2021), trapped ions Häffner et al. (2008); Monroe and Kim (2013); Bruzewicz et al. (2019); Brown et al. (2021), and neutral atom systems Wu et al. (2021); Saffman et al. (2010); Bluvstein et al. (2022, 2024) are widely adopted in QC. The superconducting circuit system is known for its fast and reliable gate operations. The trapped ions system is famous for its connectivity and long coherence while the neutral atom system is good for parallel operations and integration. Although there are other physical systems also supporting the AQM well, e.g., quantum dot and solid-state spin qubits, in the interest of conciseness, we mainly focus on the superconducting circuit system to discuss the support of the AQM from the physical stack. Detailed discussions of trapped ions and neutral atoms can be found in the Appendix.
Qubit states.
@@
Both transmon and fluxonium qubits have higher excited states Kjaergaard et al. (2020); Blais et al. (2021); Manucharyan et al. (2009); Koch et al. (2007), while the ions and atoms have other electronic levels. In transmon qubits, due to the anharmonicity provided by Josephson junctions, the transitions to higher excited states can be energetically distinguished from the transition between ground and excited states Koch et al. (2007); Manucharyan et al. (2009). Although it is still likely to populate higher excited states, especially for transmon qubits, the leakage can be reduced using quantum control methods, e.g., DRAG pulses Chow et al. (2010).
In the trapped ion and neutral atoms systems, qubits are made of individual ions and atoms that are trapped in electromagnetic/optical traps. The qubit states can be the electronic levels Bermudez et al. (2017); Pogorelov et al. (2021); Levine et al. (2018); Madjarov et al. (2020); Picken et al. (2018); Morgado and Whitlock (2021) or the hyperfine levels of the ion/atom Harty et al. (2014); Gaebler et al. (2016); Ballance et al. (2016); Wang et al. (2017, 2021); Sheng et al. (2018); Graham et al. (2019); Levine et al. (2019); Bluvstein et al. (2022). Ions and atoms also have other levels, which can be distinguished by the energy difference and by the transition rules.
@@
The universal control of a single superconducting qubit can be obtained through external microwave drives Chow et al. (2010); Koch et al. (2007); Manucharyan et al. (2009); Houck et al. (2008); Blais et al. (2021). Specifically, a superconducting qubit can be driven by microwave fields to apply arbitrary angle Pauli-X and Pauli-Y gates. The Pauli-Z rotations can be implemented virtually. The universal control over a single-qubit Hilbert space can be achieved.
@@
Two-qubit gates between superconducting qubits can also be implemented. For example, control-phase gates can be implemented by tuning the frequency of one transmon through flux drives DiCarlo et al. (2010); Barends et al. (2013), while cross-resonance gates between two fixed-frequency transmons or fluxoniums can be implemented using microwave drives Rigetti and Devoret (2010); Chow et al. (2011); Nesterov et al. (2022); Dogan et al. (2023). Furthermore, the recent development of tunable couplers enables tunable couplings between transmon qubits, which enables CZ gates with fidelity reaching $99.8\%$ with about $40$ ns Stehlik et al. (2021); Sung et al. (2021); Marxer et al. (2023). The high-fidelity two-qubit gates can also be realized. An iSWAP gate with gate time $50$ ns can reach fidelity $99.72\%$ Bao et al. (2022), while microwave-activated CZ gates can be expected to be realized with fidelity $99.9\%$ within $100$ ns Nesterov et al. (2018); Nguyen et al. (2022). With all these single-qubit and two-qubit gates, a universal gate set is supported.
@@
The universal control of the qubits made of ions or atoms can also be realized using microwave and optical methods Hilder et al. (2022); Ospelkaus et al. (2011); Harty et al. (2014); Yavuz et al. (2006); Xia et al. (2015); Bluvstein et al. (2022); Levine et al. (2022). The complete Hilbert space spaced by the qubit states are addressable. Fast and high-fidelity two-qubit gate operations are achieved Gaebler et al. (2016); Ballance et al. (2016); Saner et al. (2023); Levine et al. (2019); Bluvstein et al. (2022); Evered et al. (2023); Bluvstein et al. (2023, 2024).
@@
Although the gate operations still suffer from errors, there are active explorations on building quantum error correction codes to mitigate the imperfection Krinner et al. (2022); Acharya et al. (2023a); Bermudez et al. (2017); Bluvstein et al. (2024); Stricker et al. (2020).
Connectivity.
@@
However, the coupling between superconducting qubits is limited, where a single superconducting qubit can only couple to the qubits nearby (mostly physically connected). Therefore, the connectivity is usually limited to the nearest neighbors Blais et al. (2021).
@@
Compared to the superconducting qubits, the trapped ions and neutral atoms can maintain higher connectivity. Due to the nature of the mechanics of two-qubit gates between ions, it is possible to perform long-range two-qubit gates between ions trapped in a single trap Sørensen and Mølmer (1999). In the neutral atom system, the coherent transport of atoms enables shuttling atoms to another site to perform gate operations Bluvstein et al. (2022, 2024).
Long coherence time.
@@
Transmon and fluxonium qubits can preserve a long coherence time compared to the gate time. With the current improvement of material choices and fabrication techniques, the lifetime of transmon qubits has been improved from $\sim 1~{}\mu$Schreier et al. (2008); Houck et al. (2008) to $100$ Barends et al. (2013); Chang et al. (2013); Jin et al. (2015) to $500~{}\mu$Place et al. (2021); Wang et al. (2022a), while the coherence time $T_{2}^{*}$ can reach $0.3$ ms Place et al. (2021); Wang et al. (2022a). The coherence time can be further improved using dynamical decoupling to $0.557$ ms Wang et al. (2022a). the fluxonium qubit is less sensitive to charge noise, which extends its coherence time $100$ to $300~{}\mu$s. Recently, a fluxonium qubit with coherence time $T_{2}^{*}$ reaches $1.48$ ms has been reported Somoroff et al. (2023)
@@
Compared to superconducting qubit systems, the trapped ions and neutral atoms can have longer coherence time, especially when the qubit is encoded into the hyperfine levels of the ions/atoms. For example, the coherence time of an ion qubit can reach $5500$ s Wang et al. (2021).
Qubit readout.
@@
The state of transmon or fluxonium qubits can be read out by dispersive coupling to a microwave field Wallraff et al. (2004). The phase accumulated by the microwave field when the qubit is in its ground or excited state is different. Using the phase information of the microwave field, quantum nondemolition measurement on qubit state can be achieved in 40 ns to 100 ns with fidelity 99.0% to 99.7% Walter et al. (2017); Sunada et al. (2022); Swiadek et al. (2023); Sunada et al. (2024). Although the measurements via this method are only projective measurements along the computational basis, more general measurements, e.g, the stabilizer measurements and error corrections used in quantum error correction codes, can be realized using projective measurements on an auxiliary qubit, which is widely used in quantum error correction Krinner et al. (2022); Acharya et al. (2023b).
@@
Although superconducting qubits fully satisfy the requirements of the released AQM, they still suffer from finite gate errors. In order to reduce the gate errors, several attempts of using superconducting qubits to build error correction codes have been demonstrated Andersen et al. (2020); Krinner et al. (2022); Acharya et al. (2023a); Gong et al. (2021b). We conclude that superconducting qubits meet all the requirements from the released AQM, and have nearly all mitigation techniques available, which makes superconducting qubits one of the most promising systems for QC. But as the current construction of superconducting qubit QC systems still suffers from finite gate errors and the connectivity in a large integration of qubits, there are opportunities to use superconducting qubits as a small but fast quantum arithmetic unit in the future QC design Liu et al. (2023b).
IV.2.2 Nitrogen-Vacancy centers

### v2

@@ 4 AQM for Algorithm-Device codesign
States Operation Connectivity Coherence Readout
QA $\checkmark$ Relaxed Critical to maintain high connectivity $\checkmark$ Computational basis
QRW Encoding methods can be different. Released Based on the problem, can be relaxed. $\checkmark$ Computational basis
MBQC Physical qubits are not always alive. Relaxed (between physical qubits) Resource state generation. $\checkmark$ Arbitrary basis
AQS $\checkmark$ Relaxed Based on the problem. $\checkmark$ Computational basis (depending on problems).
QM $\checkmark$ Relaxed Connectivity to computing component is necessary. Long coherence time. Relaxed
4.1 Applications without the complete AQM
@@
4.1.1 Quantum Annealing
Analogous to the classical annealing method, quantum annealing is a QC paradigm to solve combinatorial optimization problems [52, 53, 54]. Unlike the quantum circuit computing model, quantum annealing is an analog quantum computation framework. The optimization problem is encoded into the device Hamiltonian, whose ground state corresponds to the global optimum of the problem. Solving the problem corresponds to annealing the system states from an initial state to the ground state of the problem Hamiltonian.
@@ 4.1.1 Quantum Annealing
Therefore, quantum annealing does not require the support for Rule 2 of the AQM. As the computation is not driven by quantum gate operation, there is no need to support a full universal gate set. Instead, as long as the physical device can realize the Hamiltonian $H_{0}$ and $H_{1}$ that encodes the optimization problem, e.g., the nearest-neighbor spin-spin coupling and single-qubit Pauli-X rotations [55], and can slowly tune the system Hamiltonian and readout the final state of the system, the system can support the quantum annealing algorithm to solve the specific type of optimization problems.
For example, one type of physical system commonly considered in quantum annealing is the Ising spin glass, where $H_{0}$ contains single-site $\sigma^{(x)}$ terms, while $H_{1}$ contains two-site $Z$ couplings, i.e., $\sigma^{(z)}_{i}\sigma^{(z)}_{j}$. Although the coupling terms are restricted to two-body interactions, NP-hard problems can be cast into quadratic forms using ancilla [53, 56]. A survey of NP problems mapping to spin systems can be found in Ref. [56]. Note that realizing the Hamiltonian terms does not lead to universal control of individual spin in the system.
In order to keep the resource cost low when encoding the optimization problems into a quantum annealing setup, the physical systems need to have high connectivity, which is Rule 3 of the AQM. However, there are embedding techniques that allow mapping a quantum annealing problem onto physical systems with limited connectivity [57, 58, 59].
In conclusion, the quantum annealing applications require Rule 1 and 4 of the complete AQM, while the Rule 2 and 5 can be relaxed. In addition, the Rule 3 is also critical to keep the resource cost low.
4.1.2 Quantum random walk.
Quantum random walk is a computational paradigm for QC, especially for solving graph problems. There are broad categories of quantum walk computations. In this section, we do not aim to give a thorough review of quantum random walk. For a review of quantum random walk for QC, we refer to Refs. [18, 19]. Instead, we focus on the coined discrete-time quantum walk as an example.
@@
The quantum walk model can be used to implement universal QC [20, 21]. The quantum random walk model needs a different set of requirements compared to the AQM. For example, quantum information can be encoded not into the internal state of the qubit, but instead, into the spacial modes that can hold a physical particle [60]. Other than supporting a universal gate set on qubits, quantum walk computation only requires that the physical qubit can hop between different spatial modes that encode the quantum information to carry out the computation.
In summary, quantum walk applications require Rule 4 of the AQM. However, they have distinct requirements from the general AQM. Adaptable to problem demands, the quantum walk can operate on systems with more than two quantum states by encoding the coin system into a multi-level system based on walking choices, which differs from the Rule 1 of the AQM. Rule 2 of the AQM is relaxed, as universal controllability and a universal gate set are not essential while controlled-walk operation suffices. Rule 3 can also be partially relaxed, where the unique connectivity involves linking the systems encoding coin and space degrees of freedom (DOFs). Full connectivity is not required for unreachable space DOFs. For readout, the quantum walk model primarily emphasizes distribution in space DOFs [61, 62], making complete quantum information extraction (e.g., state tomography) unnecessary. Therefore, Rule 5 is also relaxed.
4.1.3 Measurement-based quantum computation.
MBQC is a computation model different from the quantum circuit computation model [14, 15, 16, 17]. In MBQC, the computation is driven by measurements of qubits in an entangled state of qubits. This entangled state is the resource of MBQC. A two-dimensional and higher-dimensional cluster state can be used as the resource state of MBQC [15, 63, 64].
As demonstrated in Ref. [15], a set of universal gates can be equivalently performed in MBQC. Therefore, MBQC is equivalent to the quantum circuit computation model and it can support universal QC. However, compared to the requirements placed on the AQM, MBQC releases a few requirements. The physical qubits used in the MBQC are not accessible at all times, and they do not need to support a universal gate set for computation. Furthermore, the connectivity requirements are also different from the AQM, as it only requires that the physical qubits need to be able to prepare a specific resource state for MBQC. During the computation, connectivity between different physical qubits is not necessary.
@@
In summary, MBQC requires Rule 4 and 5 of the AOM. In contrast, MBQC relaxes Rule 1 of the AQM, where the physical qubits do not need to be alive throughout the computation. Rule 2 of the AQM is also relaxed, as the physical qubits do not need to perform entangling gates. Rule 3 is supported in a different manner, which refers to the connectivity of the resource state.
4.1.4 Analogue Quantum Simulation
@@
There are two main quantum simulation categories based on the type of quantum systems used. One involves digital quantum computers equipped with algorithms to compute physical properties, such as the VQE algorithm designed for determining molecular ground states and energies [47]. The other method focuses on employing quantum devices for analogue simulations [65, 66, 67, 68], which is the emphasis of this subsection.
@@
Analogue quantum simulations can address low-energy steady states or dynamics of physical systems. Solving the ground state of a physical system can use adiabatic evolution, akin to the discussion of the quantum annealing method (see Sec. 4.1.1). Alternatively, when the problem directly maps to a quantum system, i.e., the quantum simulator can realize the target Hamiltonian, manipulating the simulator can simulate the dynamics of the target system. Therefore, the measurement of the quantum simulator can help to reveal the physical properties of the target problem. For example, the Ising model can naturally map to a neutral atom lattice [67]. The quantum many-body phase transition can be directly observed by manipulating the atom interactions [67, 69].
@@
In summary, analogue quantum simulation still demands Rule 1 and 4 of the AQM. As the simulator does not need to support a universal gate set for computing, or achieving universal control on each qubit, as long as the target Hamiltonian can be realized, Rule 2 of the AQM is relaxed. As the connectivity and the readout capability depend on the physical problems being simulated, Rule 3 and 5 of the AQM have the potential to be relaxed.
4.1.5 Quantum Memory
@@
Quantum memory is an application of quantum systems, which can preserve quantum information for an extended period of time. Unlike the other quantum computation paradigms and quantum simulation schemes, quantum memory does not aim to perform computation. Therefore, the quantum memory application has different demands on the AQM [51].
@@
To support quantum memory applications, Rule 1 and 4 of the AQM are required. The AQM Rule 2 can be relaxed, as a quantum system does not need to support universal control of individual qubits and a universal gate set for quantum computation. Rule 5 can also be relaxed as when the quantum memory acts as a computing module in the future QPU architecture design, its readout capability is not necessary. Rule 3 is differently demanded, as the connectivity of the qubits inside the quantum memory modules is not essential, while connectivity to computing qubits is necessary [51].
4.2 Physical systems supporting the AQM
@@
In this section, we examine the physical system from the lower stacks of the AQM. We do not aim to provide a thorough review of useful physical systems for QC and QIP. Instead, we mainly focus on three types of physical systems: (1) physical systems with strong support of the AQM, e.g., superconducting circuits, trapped ions, and neutral atoms, (2) systems with unique strength and weakness in supporting the AQM, e.g., nitrogen-vacancy centers in diamond crystals, and photonic systems, (3) systems with partial support of the AQM, e.g, the quantum memory systems for photons.
4.2.1 Physical systems for AQM support
@@
Superconducting circuits [70, 71, 72, 73, 74], trapped ions [75, 76, 77, 78], and neutral atom systems [79, 80, 81, 12] are widely adopted in QC. The superconducting circuit system is known for its fast and reliable gate operations. The trapped ions system is famous for its connectivity and long coherence while the neutral atom system is good for parallel operations and integration. Although there are other physical systems that also support the AQM well, e.g., quantum dot and solid-state spin qubits, in the interest of conciseness, we mainly focus on the superconducting circuit system to discuss the support of the AQM from the physical stack. Detailed discussions of trapped ions and neutral atoms can be found in the Appendix.
a. Qubit states.
@@
Both transmon and fluxonium qubits have higher excited states [72, 74, 82, 83], while the ions and atoms have other electronic levels. In transmon qubits, due to the anharmonicity provided by Josephson junctions, the transitions to higher excited states can be energetically distinguished from the transition between ground and excited states [83, 82]. Although it is still likely to populate higher excited states, especially for transmon qubits, the leakage can be reduced using quantum control methods, e.g., DRAG pulses [84].
In the trapped ion and neutral atoms systems, qubits are made of individual ions and atoms that are trapped in electromagnetic/optical traps. The qubit states can be the electronic levels [85, 86, 87, 88, 89, 90] or the hyperfine levels of the ion/atom [91, 92, 93, 94, 95, 96, 97, 98, 81]. Ions and atoms also have other levels, which can be distinguished by the energy difference and by the transition rules.
@@
The universal control of a single superconducting qubit can be obtained through external microwave drives [84, 83, 82, 99, 74]. Specifically, a superconducting qubit can be driven by microwave fields to apply arbitrary angle Pauli-X and Pauli-Y gates. The Pauli-Z rotations can be implemented virtually. The universal control over a single-qubit Hilbert space can be achieved.
@@
Two-qubit gates between superconducting qubits can also be implemented. For example, control-phase gates can be implemented by tuning the frequency of one transmon through flux drives [100, 101], while cross-resonance gates between two fixed-frequency transmons or fluxoniums can be implemented using microwave drives [102, 103, 104, 105]. Furthermore, the recent development of tunable couplers enables tunable couplings between transmon qubits, which enables CZ gates with fidelity reaching $99.8\%$ with about $40$ ns [106, 107, 108]. The high-fidelity two-qubit gates can also be realized. An iSWAP gate with gate time $50$ ns can reach fidelity $99.72\%$ [109], while microwave-activated CZ gates can be expected to be realized with fidelity $99.9\%$ within $100$ ns [110, 111]. With all these single-qubit and two-qubit gates, a universal gate set is supported.
@@
The universal control of the qubits made of ions or atoms can also be realized using microwave and optical methods [112, 113, 91, 114, 115, 81, 116]. The complete Hilbert space spanned by the qubit states is addressable. Fast and high-fidelity two-qubit gate operations are achieved [92, 93, 117, 98, 81, 118, 119, 12].
@@
Although the gate operations still suffer from errors, there are active explorations on building quantum error correction codes to mitigate the imperfection [120, 121, 85, 12, 122].
c. Connectivity.
@@
However, the coupling between superconducting qubits is limited, where a single superconducting qubit can only couple to the qubits nearby (mostly physically connected). Therefore, the connectivity is usually limited to the nearest neighbors [74].
@@
Compared to superconducting qubits, the trapped ions and neutral atoms can maintain higher connectivity. Due to the nature of the mechanics of two-qubit gates between ions, it is possible to perform long-range two-qubit gates between ions trapped in a single trap [123]. In the neutral atom system, the coherent transport of atoms enables shuttling atoms to another site to perform gate operations [81, 12].
d. Long coherence time.
@@
Transmon and fluxonium qubits can preserve a relatively long coherence time compared to the gate time. With the current improvement of material choices and fabrication techniques, the lifetime of transmon qubits has been improved from $\sim 1\leavevmode\nobreak\ \mu$[124, 99] to $100$ [101, 125, 126] to $500\leavevmode\nobreak\ \mu$[127, 128], while the coherence time $T_{2}^{*}$ can reach $0.3$ ms [127, 128]. The coherence time can be further improved using dynamical decoupling to $0.557$ ms [128]. the fluxonium qubit is less sensitive to charge noise, which extends its coherence time $100$ to $300\leavevmode\nobreak\ \mu$s. Recently, a fluxonium qubit with coherence time $T_{2}^{*}$ reaches $1.48$ ms has been reported [129]
@@
Compared to superconducting qubit systems, trapped ions and neutral atoms can have longer coherence time, especially when the qubit is encoded into the hyperfine levels of the ions/atoms. For example, the coherence time of an ion qubit can reach $5500$ s [95].
e. Qubit readout.
@@
The state of transmon or fluxonium qubits can be read out by dispersive coupling to a microwave field [130]. The phase accumulated by the microwave field when the qubit is in its ground or excited state is different. Using the phase information of the microwave field, quantum nondemolition measurement on qubit state can be achieved in 40 ns to 100 ns with fidelity 99.0% to 99.7% [131, 132, 133, 134]. Although the measurements via this method are only projective measurements along the computational basis, more general measurements, e.g, the stabilizer measurements and error corrections used in quantum error correction codes, can be realized using projective measurements on an auxiliary qubit, which is widely used in quantum error correction [120, 135].
@@
Although superconducting qubits fully satisfy the requirements of the relaxed AQM, they still suffer from finite gate errors. In order to reduce the gate errors, several attempts of using superconducting qubits to build error correction codes have been demonstrated [136, 120, 121, 137]. We conclude that superconducting qubits meet all the requirements from the relaxed AQM, and have nearly all mitigation techniques available, which makes superconducting qubits one of the most promising systems for QC. But as the current construction of superconducting qubit QC systems still suffers from finite gate errors and the connectivity in a large integration of qubits, there are opportunities to use superconducting qubits as a small but fast quantum arithmetic unit in future QC design [51].
4.2.2 Nitrogen-Vacancy centers